OpenStudy (samara8954):

Which is an equation for the given word problem? After Nancy deposits her birthday gift of $20.00 into her savings account, her balance is $225.50. Which equation can be used to find the amount, n, in her account before she made the deposit? A. n – 225.50 = 20.00 B. 20.00 – n = 225.50 C. n + 225.50 = 20.00 D. 20.00 + n = 225.50

OpenStudy (samara8954):

@aaliyahfairgood

OpenStudy (samara8954):

@dtan5457 @jeffyblood

OpenStudy (danjs):

starting amount n add on to that the 20 dollar deposit that should now equal total in the account after the deposit

OpenStudy (samara8954):

so c

OpenStudy (samara8954):

@DanJS

OpenStudy (samara8954):

@mathstudent55

OpenStudy (aaliyahfairgood):

Before the deposit, the balance was $205.50 so it would be set up like this 20-n=225.50

OpenStudy (danjs):

C ) is described like this - she deposited 225 into her account, now the total balance = 20 dollars

OpenStudy (danjs):

she must have started in the negative value for choice C

OpenStudy (danjs):

(Starting amt n) + (deposit 20) = (total 225.50)

OpenStudy (danjs):

Starts with n, adds 20, now has 225.5 20 + n = 225.5

OpenStudy (aaliyahfairgood):

So it is not c or b, it's d